# ELIP: Enhanced Visual-Language Foundation Models for Image Retrieval

### v1

@@ B.3 Attention Map of CLIP and ELIP-C
Figure 17 shows the cross-attention map of [CLS] token on patch tokens for both CLIP and our ELIP-C on COCO (Rows 1-2) and Flickr (Rows 3-4). We display the cases where the image matches the text query on the left, and the images that do not match the text query on the right. The visualisation is based on the codebase of [10], where a cooler color represents a higher activation.
@@
It can be observed that when the image matches the text query, our generated visual prompt vector can effectively boost the selection of image features relevant to the text query: For example, when the text query is ‘A person wearing a banana headdress and necklace’ (Row 1), comparing the attention map of CLIP and ELIP-C, we can observe ELIP-C enables more attention on the banana headdress and necklace; when the text query is ‘A cute cat laying down in a sink’ (Row 2), we note that ELIP-C gives significantly more attention on the cat; when the text query is ‘A black and white dog is running in a grassy garden surrounded by a white fence’ (Row 3), it can be noted that ELIP-C largely increases the attention on the black and white dog and the white fence; when the text query is ‘A group of young people are listening to a man in a blue shirt telling them about Brazil’ (Row 4), it can be noted that ELIP-C makes the attention to focus more on the man in blue shirt who is talking, as well as the Brazil flag and people listening.
The difference is not significant if the image does not match the query (Column 4-6).
Figure 17: Visualisation of attention map of CLIP and ELIP-C shows the cross-attention map of [CLS] token on patch tokens for both CLIP and our ELIP-C on COCO and Flickr. Left: image matches the text query; Right: image does not match the text query. COCO: Rows 1-2; Flickr: Rows 3-4. For the situation where the image matches the text query, the difference in the attention maps shows our generated visual prompt vector effectively improves the selection of image features relevant to the text query. Please see text for a more detailed discussion. The visualisation is based on the codebase of [10], where a cooler color represents a higher activation.
@@
Figure 18 shows the cross-attention map of [CLS] token on patch tokens for both CLIP and our ELIP-C on Occluded COCO (Rows 1-2) and ImageNet-R (Rows 3-4). We display the cases where the image matches the text query on the left, and the images that do not match the text query on the right. The visualisation is based on the codebase of [10], where a cooler color represents a higher activation.
@@
It can be observed that when the image matches the text query, our generated visual prompt vector can effectively boost the selection of image features relevant to the text query: For example, when the text query is ‘banana’ (Row 1), it can be observed that ELIP-C enables more concentration of the attention on the banana rather than other areas; when the text query is ‘baseball bat’ (Row 2), we observe that ELIP-C brings more attention on the baseball bat; when the text query is ‘baboon’ (Row 3) and ‘ant’ (Row 4), ELIP-C makes the attention to focus more on the baboon and ant respectively.
The difference is not significant if the image does not match the query (Column 4-6).
@@
Figure 18: Visualisation of attention map of CLIP and ELIP-C shows the cross-attention map of [CLS] token on patch tokens for both CLIP and our ELIP-C on Occluded COCO and ImageNet-R. Left: image matches the text query; Right: image does not match the text query. Occluded COCO: Rows 1-2; ImageNet-R: Rows 3-4. For the situation where the image matches the text query, the difference in the attention maps shows our generated visual prompt vector effectively improves the selection of image features relevant to the text query. Please see text for a more detailed discussion. The visualisation is based on the codebase of [10], where a cooler color represents a higher activation.
B.4 Attention Map of BLIP and ELIP-B
@@
It can be observed that when the image matches the text query, our generated visual prompt vector can effectively boost the selection of image features relevant to the text query: For example, when the text query is ‘A baseball player holding a bat while standing in a field’ (Row 1), comparing the attention map of BLIP-2 and ELIP-B we can observe ELIP-B enables more attention on the baseball bat; when the text query is ‘A brown and white dog wearing a neck tie’ (Row 2), we note that ELIP-B gives more attention on the tie and the dog; when the text query is ‘A bunch of beer pull tabs at a bar with Christmas lights on the ceiling’ (Row 3), it can be noted that ELIP-B largely increases the attention on the Christmas lights on the ceiling and the bunch of beer pull tabs; when the text query is ‘A father and daughter holding a young tree upright, ready to be planted, as the son stands to their side wielding a shovel’ (Row 4), it can be noted that ELIP-B makes the attention to focus more on the father, daughter, son and tree.
The difference is not significant if the image does not match the query (Column 4-6).
Figure 19: Visualisation of attention map of BLIP-2 and ELIP-B shows the cross-attention map of the query tokens at the ITM head on patch tokens for both BLIP-2 and our ELIP-B on COCO and Flickr. There are 32 query tokens in total, and we take the average value of the 32 query tokens to visualise the attention map. The attention map represents image features at which locations are selected and concentrated. Left: image matches the text query; Right: image does not match the text query. COCO: Rows 1-2; Flickr: Rows 3-4. For the situation where the image matches the text query, the difference in the attention maps shows our generated visual prompt vector effectively improves the selection of image features relevant to the text query. Please see text for a more detailed discussion. The visualisation is based on the codebase of [68], where a warmer color represents a higher activation.
@@
Figure 20 shows the cross-attention map of the query tokens at the ITM head on patch tokens for both BLIP-2 and our ELIP-B on Occluded COCO (Rows 1-2) and ImageNet-R (Rows 3-4). There are 32 query tokens in total, and we take the average value of the 32 query tokens to visualise the attention map. The attention map represents image features at which locations are selected and concentrated. We display the situations where the image matches the text query on the left, and where the images do not match the text query on the right. The visualisation is based on the codebase of [68], where a warmer color represents a higher activation.
It can be observed that when the image matches the text query, our generated visual prompt vector can effectively boost the selection of image features relevant to the text query: For example, when the text query is ‘surfboard’ (Row 1), it can be observed that ELIP-B enables more concentration of the attention on the surfboard; when the text query is ‘tennis racket’ (Row 2), we observe that ELIP-B brings more attention on the tennis racket; when the text query is ‘backpack’ (Row 3) and ‘basketball’ (Row 4), ELIP-B makes the attention to focus more on the backpack and basketball respectively.
The difference is not significant if the image does not match the query (Column 4-6).
@@
Figure 20: Visualisation of attention map of BLIP-2 and ELIP-B shows the cross-attention map of the query tokens at the ITM head on patch tokens for both BLIP-2 and our ELIP-B on Occluded COCO and ImageNet-R. There are 32 query tokens in total, and we take the average value of the 32 query tokens to visualise the attention map. The attention map represents image features at which locations are selected and concentrated. Left: image matches the text query; Right: image does not match the text query. Occluded COCO: Rows 1-2; ImageNet-R: Rows 3-4. For the situation where the image matches the text query, the difference in the attention maps shows our generated visual prompt vector effectively improves the selection of image features relevant to the text query. Please see text for a more detailed discussion. The visualisation is based on the codebase of [68], where a warmer color represents a higher activation.
@@ Appendix C Ablation Study for Different Number of Generated Visual Prompt Tokens
Table 5: Ablation study on number of generated visual prompt tokens. Following Table 1, we use CLIP as the baseline for the ablation study. 0 refers to the CLIP baseline and a non-zero number refers to our ELIP-C architecture with different number of generated visual prompt vectors.
As mentioned in Section 7.1, here we study the performance as the number of generated visual prompt tokens are varied. The results are shown in Table 5, and it can be observed that generating multiple tokens, e.g., 2 tokens (Setting $\mathbb{D}$), 5 tokens (Setting $\mathbb{D}$) or 10 tokens (Setting $\mathbb{E}$), brings an improvement compared with generating only 1 token (Setting $\mathbb{A}$), and Setting $\mathbb{E}$ performs best. Therefore, we generate 10 tokens in all experiments.

### v2

@@ B.4 Attention Map of CLIP and ELIP-C
Figure 21 shows the cross-attention map of [CLS] token on patch tokens for both CLIP and our ELIP-C on COCO (Rows 1-2) and Flickr (Rows 3-4). We display the cases where the image matches the text query on the left and the images that do not match the text query on the right. The visualisation is based on the codebase of [71], where a warmer color represents a higher activation.
@@
Figure 21: Visualisation of attention map of CLIP and ELIP-C shows the cross-attention map of [CLS] token on patch tokens for both CLIP and our ELIP-C on COCO and Flickr. Left: image matches the text query; Right: image does not match the text query. COCO: Rows 1-2; Flickr: Rows 3-4. It can be observed that when the image matches the text query, our generated visual prompt vectors can effectively boost the selection of image features relevant to the text query: For example, when the text query is ‘A person wearing a banana headdress and necklace’ (Row 1), comparing the attention map of CLIP and ELIP-C, we can observe ELIP-C enables more attention on the banana headdress and necklace; when the text query is ‘A cute cat laying down in a sink’ (Row 2), we note that ELIP-C gives significantly more attention on the cat; when the text query is ‘A black and white dog is running in a grassy garden surrounded by a white fence’ (Row 3), it can be noted that ELIP-C largely increases the attention on the black and white dog and the white fence; when the text query is ‘A group of young people are listening to a man in a blue shirt telling them about Brazil’ (Row 4), it can be noted that ELIP-C makes the attention to focus more on the man in a blue shirt who is talking, as well as the Brazil flag and people listening. The difference is not significant if the image does not match the query (Columns 4-6).
@@
Figure 22 shows the cross-attention map of [CLS] token on patch tokens for both CLIP and our ELIP-C on Occluded COCO (Rows 1-2) and ImageNet-R (Rows 3-4). We display the cases where the image matches the text query on the left and the images that do not match the text query on the right. The visualisation is based on the codebase of [71], where a warmer color represents a higher activation.
@@
Figure 22: Visualisation of attention map of CLIP and ELIP-C shows the cross-attention map of [CLS] token on patch tokens for both CLIP and our ELIP-C on Occluded COCO and ImageNet-R. Left: image matches the text query; Right: image does not match the text query. Occluded COCO: Rows 1-2; ImageNet-R: Rows 3-4. It can be observed that when the image matches the text query, our generated visual prompt vectors can effectively boost the selection of image features relevant to the text query: For example, when the text query is ‘banana’ (Row 1), it can be observed that ELIP-C enables more concentration of attention on the banana rather than other areas; when the text query is ‘baseball bat’ (Row 2), we observe that ELIP-C brings more attention on the baseball bat; when the text query is ‘baboon’ (Row 3) and ‘ant’ (Row 4), ELIP-C makes the attention to focus more on the baboon and ant respectively. The difference is not significant if the image does not match the query (Columns 4-6).
B.5 Attention Map of SigLIP/SigLIP-2 and ELIP-S/ELIP-S-2
@@
Figure 23 shows the cross-attention map of [CLS] token on patch tokens for both SigLIP/SigLIP-2 and our ELIP-S/ELIP-S-2 on COCO (Rows 1 and 3) and Flickr (Rows 2 and 4). We display the cases where the image matches the text query on the left and the images that do not match the text query on the right. The visualisation is based on the codebase of [71], where a warmer color represents a higher activation.
Figure 23: Visualisation of attention map of SigLIP/SigLIP-2 and ELIP-S/ELIP-S-2 shows the cross-attention map of [CLS] token on patch tokens for both SigLIP/SigLIP-2 and our ELIP-S/ELIP-S-2 on COCO and Flickr. Top: SigLIP v.s. ELIP-S; Bottom: SigLIP-2 v.s. ELIP-S-2. Left: image matches the text query; Right: image does not match the text query. COCO: Rows 1 and 3; Flickr: Rows 2 and 4. It can be observed that when the image matches the text query, our generated visual prompt vectors can effectively boost the selection of image features relevant to the text query: For example, when the text query is ‘A basket ball player is posing in front of a basket’ (Row 1), comparing the attention map of SigLIP and ELIP-S, we can observe ELIP-S enables more attention on the basketball, the basket, and the player; when the text query is ‘A black dog is slowly crossing a fallen log that is outstretched over a stream of water’ (Row 2), we note that ELIP-S gives significantly more attention on the black dog and the log; when the text query is ‘A dark skinned child getting ready to be pushed on a swing’ (Row 3), it can be noted that ELIP-S-2 largely increases the attention on the child and the swing; when the text query is ‘A boy is hanging out of the window of a yellow taxi’ (Row 4), it can be noted that ELIP-S-2 makes the attention to focus more on the boy out of the window. The difference is not significant if the image does not match the query (Columns 4-6).
Figure 24 shows the cross-attention map of [CLS] token on patch tokens for both SigLIP/SigLIP-2 and our ELIP-S/ELIP-S-2 on Occluded COCO (Rows 1 and 3) and ImageNet-R (Rows 2 and 4). We display the cases where the image matches the text query on the left and the images that do not match the text query on the right. The visualisation is based on the codebase of [71], where a warmer color represents a higher activation.
Figure 24: Visualisation of attention map of SigLIP/SigLIP-2 and ELIP-S/ELIP-S-2 shows the cross-attention map of [CLS] token on patch tokens for both SigLIP/SigLIP-2 and our ELIP-S/ELIP-S-2 on Occluded COCO and ImageNet-R. Top: SigLIP v.s. ELIP-S; Bottom: SigLIP-2 v.s. ELIP-S-2. Left: image matches the text query; Right: image does not match the text query. Occluded COCO: Rows 1 and 3; ImageNet-R: Rows 2 and 4. It can be observed that when the image matches the text query, our generated visual prompt vectors can effectively boost the selection of image features relevant to the text query: For example, when the text query is ‘bird’ (Row 1), it can be observed that ELIP-S enables more concentration of attention on the bird; when the text query is ‘beagle’ (Row 2), we observe that ELIP-S brings more attention on the beagle; when the text query is ‘book’ (Row 3) and ‘chow chow’ (Row 4), ELIP-S-2 makes the attention to focus more on the book and chow chow respectively. The difference is not significant if the image does not match the query (Columns 4-6).
B.6 Attention Map of BLIP-2 and ELIP-B
@@
Figure 25: Visualisation of attention map of BLIP-2 and ELIP-B shows the cross-attention map of the query tokens at the ITM head on patch tokens for both BLIP-2 and our ELIP-B on COCO and Flickr. There are 32 query tokens in total, and we take the average value of the 32 query tokens to visualise the attention map. The attention map represents image features at which locations are selected and concentrated. Left: image matches the text query; Right: image does not match the text query. COCO: Rows 1-2; Flickr: Rows 3-4. It can be observed that when the image matches the text query, our generated visual prompt vectors can effectively boost the selection of image features relevant to the text query: For example, when the text query is ‘A baseball player holding a bat while standing in a field’ (Row 1), comparing the attention map of BLIP-2 and ELIP-B we can observe ELIP-B enables more attention on the baseball bat; when the text query is ‘A brown and white dog wearing a neck tie’ (Row 2), we note that ELIP-B gives more attention on the tie and the dog; when the text query is ‘A bunch of beer pull tabs at a bar with Christmas lights on the ceiling’ (Row 3), it can be noted that ELIP-B largely increases the attention on the Christmas lights on the ceiling and the bunch of beer pull tabs; when the text query is ‘A father and daughter holding a young tree upright, ready to be planted, as the son stands to their side wielding a shovel’ (Row 4), it can be noted that ELIP-B makes the attention focus more on the father, daughter, son and tree. The difference is not significant if the image does not match the query (Columns 4-6).
@@
Figure 26 shows the cross-attention map of the query tokens at the ITM head on patch tokens for both BLIP-2 and our ELIP-B on Occluded COCO (Rows 1-2) and ImageNet-R (Rows 3-4). There are 32 query tokens in total, and we take the average value of the 32 query tokens to visualise the attention map. The attention map represents image features at which locations are selected and concentrated. We display the situations where the image matches the text query on the left and where the images do not match the text query on the right. The visualisation is based on the codebase of [71], where a warmer color represents a higher activation.
@@
Figure 26: Visualisation of attention map of BLIP-2 and ELIP-B shows the cross-attention map of the query tokens at the ITM head on patch tokens for both BLIP-2 and our ELIP-B on Occluded COCO and ImageNet-R. There are 32 query tokens in total, and we take the average value of the 32 query tokens to visualise the attention map. The attention map represents image features at which locations are selected and concentrated. Left: image matches the text query; Right: image does not match the text query. Occluded COCO: Rows 1-2; ImageNet-R: Rows 3-4. It can be observed that when the image matches the text query, our generated visual prompt vectors can effectively boost the selection of image features relevant to the text query: For example, when the text query is ‘pizza’ (Row 1), it can be observed that ELIP-B enables more concentration of attention on the pizza; when the text query is ‘surfboard’ (Row 2), we observe that ELIP-B brings more attention on the surfboard; when the text query is ‘bee’ (Row 3) and ‘basketball’ (Row 4), ELIP-B makes the attention to focus more on the bee and basketball respectively. The difference is not significant if the image does not match the query (Columns 4-6).
@@ Appendix C Ablation Study for Different Number of Generated Visual Prompt Tokens
Table 5: Ablation study on number of generated visual prompt tokens. Following Table 1 of the main paper, we use CLIP as the baseline for the ablation study. 0 refers to the CLIP baseline and a non-zero number refers to our ELIP-C architecture with different numbers of generated visual prompt vectors.
As mentioned in Section 7.1, here we study the performance as the number of generated visual prompt tokens are varied. The results are shown in Table 5, and it can be observed that generating multiple tokens, e.g., 2 tokens (Setting $\mathbb{C}$), 5 tokens (Setting $\mathbb{D}$) or 10 tokens (Setting $\mathbb{E}$), brings an improvement compared with generating only 1 token (Setting $\mathbb{B}$), and Setting $\mathbb{E}$ performs best. Therefore, we generate 10 tokens in all experiments.
Appendix D Further Ablation Study for ELIP-B
Settings Fine-tune ITM Head JEST Selection COCO Flickr
Recall@1 Recall@5 Recall@10 Recall@1 Recall@5 Recall@10
$\mathbb{A}$ 68.25 87.72 92.63 89.74 98.18 98.94
$\mathbb{B}$ 68.26 87.86 92.78 89.80 98.26 99.14
$\mathbb{C}$ 67.35 87.38 92.53 89.08 98.06 99.08
$\mathbb{D}$ 68.41 87.88 92.78 90.08 98.34 99.22
Table 6: Ablation study of ELIP-B in terms of training data selection inspired by JEST [25] and fine-tuning ITM head. Setting $\mathbb{A}$ is the baseline BLIP-2. Setting $\mathbb{B}$ is trained on a random subset of the same size of Settings $\mathbb{C}$ and $\mathbb{D}$. It can be observed that both the strategies of fine-tuning the ITM head and training on selected subset by JEST contribute to the improvement of our ELIP-B.
In this section, we have further conducted an ablation study on ELIP-B to show the effectiveness of selecting the training data using the strategy inspired by JEST [25] as described in Section 5.2, as well as also fine-tuning the ITM head as in Figure 3 of the main paper, as these are two unique things we only applied to BLIP-2.
In Table 6, we can observe that: 1) Comparing Settings $\mathbb{B}$ and $\mathbb{D}$, we can note the effectiveness of training on JEST selected training batches compared with selecting batches randomly; 2) Comparing Settings $\mathbb{C}$ and $\mathbb{D}$, we can infer that it is important to also fine-tune the lightweight ITM head as the input features to the ITM head have been changed in ELIP-B.
Appendix E Efficiency of ELIP Pre-Training
      Model       GPU Hours       #GPU       BS       FLOPS       $\Delta$FLOPS
      CLIP       10736 A100       176 $\times$ A100       33792       33.7G       -
      ELIP-C       144 A40       2 $\times$ A40       40       35.8G       2.1G
      SigLIP       1152 TPUv4       16 $\times$ TPUv4       32768       604.5G       -
      ELIP-S       144 A40       2 $\times$ A40       10       613.3G       8.8G
      SigLIP-2       *       2048 $\times$ TPUv5e       32768       1.31T       -
      ELIP-S-2       144 A40       2 $\times$ A40       10       1.34T       0.03T
      BLIP-2       2304 A100       16 $\times$ A100       1680       1.33T       -
      ELIP-B       48 A40       2 $\times$ A40       12       1.35T       0.02T
Table 7: Efficiency of ELIP pre-training compared with original pre-training. Our method significantly improves the efficiency on training time, GPU requirement, and batch size, introducing only a marginal increase on FLOPS for our trainable MLP mapping network. *SigLIP-2 original training time is not provided in their paper.
As mentioned in Section 7, in the table below we show a comparison of our method to the original pre-training of CLIP, SigLIP, SigLIP-2, and BLIP-2 in terms of training time, GPU requirement, batch size, and FLOPS. It can be observed that our method significantly improves the efficiency on training time, GPU requirement, and batch size, while only introducing a marginal increase on FLOPS which results from our trainable MLP mapping network.